# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 3.2 Self-improving AI algorithm
    Evaluation: The system outcome $\Phi(\Pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\Pi_{t},c_{t},s_{t})\}$.
3.3 Bayesian Optimization (BO)
As discussed in Section 1, we would like to formulate self-improving AI as parallel gradient-based UCB-BO. At its core, BO maximizes an unknown and expensive-to-evaluate objective function under a limited evaluation budget. Formally, consider a black-box function $f:\mathcal{Z}\to\mathbb{R}$, where in our setting $f$ corresponds to the system-level objective $J(\Pi)$ over prompts. BO maintains a surrogate model, typically a Gaussian Process (GP) or another probabilistic regressor defined on $\mathcal{Z}$, that summarizes current beliefs about $f$. For any candidate $z\in\mathcal{Z}$, the surrogate yields a posterior mean $\mu(z)$ and an associated uncertainty measure $\sigma(z)$.
In BO, the following evaluation point is selected by choosing the point that maximizes an acquisition function $A(z)$ constructed from $\mu(z)$ and $\sigma(z)$, rather than by directly maximizing $\mu(z)$. Acquisition functions encode an explicit trade-off between exploitation (preferring points with high predicted value) and exploration (preferring points with high uncertainty). A widely used choice is the Upper Confidence Bound (UCB) acquisition function
$A_{\beta}(z)\;=\;\mu(z)\;+\;\beta\,\sigma(z),$ (6)
where the exploration parameter $\beta>0$ controls the strength of the exploration term (srinivas2012information). Intuitively, UCB acquisition function gives a positive bonus to unexplored areas in addition to the posterior mean. This allows us to follow the “optimism in the face of uncertainty” principle, which is known to trade off exploration and exploitation effectively. Indeed, UCB-BO is an optimal algorithm for achieving evaluation efficiency, in the sense that no algorithm can achieve better regret guarantees (wang2023regret; whitehouse2023sublinear) (See Web Appendix E for a detailed theoretical discussion of BO.)
In practice, since the maxima of the UCB acquisition function are often not available in a closed form, we typically use parallel gradient-based BO with the UCB acquisition function (wilson2018maximizing). In parallel gradient-based BO, we start from multiple initial points and take gradients from them to climb up the UCB acquisition function to find its local maxima. We then choose the points to evaluate among the local maxima (Figure 3).
In our setting, we do not construct an explicit $\mu(\cdot)$ and $\sigma(\cdot)$ (and their corresponding surrogate model and acquisition function) for the prompt space; instead, we interpret $\mu(\cdot)$ and $\sigma(\cdot)$ as implicit beliefs within the LLM and design our method to emulate parallel gradient-based UCB in this implicit space. Textual gradient (textgrad) is designed to elicit gradients for $\mu(\cdot)$ in the prompt space without explicitly modeling those beliefs within the LLM. However, what we want is not gradients for $\mu(\cdot)$; we want gradients for $A_{\beta}(z)$, the UCB acquisition function.
@@ 4 Best-of-N gradient and Theory
As discussed in Section 1, a core challenge in casting self-improving AI as parallel, gradient-based UCB Bayesian optimization in the space of prompts is the absence of an explicit acquisition function (e.g., UCB) and its gradient. As we discussed earlier, textual gradients (textgrad) can guide the prompt toward the highest expected improvement based on immediate feedback, as a textual gradient elicits the gradient of the posterior mean $\nabla\mu(x)$. However, following the posterior mean’s gradient rather than UCB’s is a purely exploitative strategy that often leads to suboptimal local maxima.
In this section, we propose that a gradient direction we call Best-of-N gradient, which generates $N$ stochastic textual gradients and selecting the most promising candidate via a Best-of-$N$ mechanism, can elicit the gradient of the UCB acquisition function within LLM’s mind without explicitly constructing the LLM’s implicit posterior mean and the associated uncertainty.
To operationalize the Best-of-N gradient, we should first formalize the requisite capabilities of the external Large Language Model, which we designate as the Critic LLM.
4.1 Critic LLM.
Let $\Pi$ denote a prompt and $c=\Phi(\Pi)$ denote the corresponding system outcome. We define the history of optimization up to iteration $t$ as $H_{t}=\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau\leq t}$. Then we define the Critic LLM, $M_{\text{critic}}$, as an LLM that can conduct the operations of the following four distinct operators:
@@
    Pairwise Judge Operator: A function that compares two outcome $c_{1}$ and $c_{2}$, given a reflection $R$ and predicts which outcome will yield a higher true score:
    $\textsc{PairwiseJudge}\big(c_{1},c_{2};R,M_{\text{critic}}\big)\in\{c_{1},c_{2}\}$ (7)
@@
    Reflection Operator: A function that compresses the entire history into a natural-language summary of rules or guidelines:
    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ (8)
@@
    Textual Gradient Operator: A stochastic function that generates a textual gradient $\delta$ (e.g., “Make the lighting warmer”) intended to improve the prompt:
    $\delta\sim\nabla_{\text{text}}(\Pi;R)$ (9)
@@
    Rewrite Operator: A deterministic function that maps a prompt $\Pi$ and an edit $\delta$ to a new prompt $\Pi^{\prime}$:
    $\Pi^{\prime}=\operatorname{Apply}(\Pi,\delta;M_{\text{critic}})$ (10)
In the digital-marketing example, these correspond to an LLM’s ability to (1) compare two ad creatives and predict which will perform better (calderon2025alternative), (2) examine past performance data and derive guidelines for effective ads in context (jiang2024many; ferber2024context), (3) propose prompt improvements based on those guidelines (textgrad; li2025test), and (4) apply prompt improvements to get the new prompt.
@@ 4.2 Best-of-N gradient
The notion of a Best-of-$N$ gradient builds on the widely used Best-of-$N$ strategy for test-time alignment, where an LLM trades additional inference-time (test-time) compute for higher-quality outputs by sampling multiple candidate generations and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical). In the standard Best-of-$N$ strategy for prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one. While effective for choosing a single output, our goal is to improve the prompt itself to ensure sustained performance improvement. To achieve this, we introduce the concept of the Best-of-N Gradient. Instead of sampling $N$ candidate prompts directly, we sample $N$ distinct textual gradients (directions of improvement). By applying these $N$ textual edits to the current prompt and selecting the LLM-as-a-judge-predicted most promising candidate solution (Best-of-N), we aim to identify the textual gradient direction for the implicit acquisition function. The procedure operates as follows:
@@
    Textual gradient generation: Using $M_{\text{critic}}$, query the textual-gradient operator $N$ independent times to obtain a set of small edits:
    $\delta^{(i)}\sim\nabla_{\text{text}}(\Pi;R,M_{\text{critic}}),\quad i=1,\dots,N.$ (11)
@@
    Candidate prompt generation: Apply each edit to the current prompt to generate $N$ candidate prompts:
    $\Pi^{(i)}=\operatorname{Apply}(\Pi,\delta^{(i)};M_{\text{critic}}),\quad i=1,\dots,N.$ (12)
@@
    Candidate system outcome generation: Generate the corresponding system outcomes for each candidate prompt:
    $c^{(i)}=\Phi(\Pi^{(i)}),\quad i=1,\dots,N.$ (13)
@@
    Best-of-N selection: Select the best candidate by iteratively applyng PairwiseJudge operation using $M_{\text{critic}}$ as LLM-as-a-judge. We define this selection as Best-of-N operation, which induces the index $i^{\star}$ as the best one among N:
    $i^{\star}=\textsc{Best-of-N}\big(\{(\Pi^{(i)},c^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (14)
    The resulting prompt $\Pi^{(i^{\star})}$ represents the update step taken by the Best-of-N Gradient.
For the Best-of-N selection, we employ the pairwise tournament (liu2025pairjudgermperformbestofn). First, the $N$ candidates are permuted to produce randomly chosen tournament matches. Each match compares two outcomes by querying $M_{\text{critic}}$ for a head-to-head judgment (For the prompt, please refer to Figure 11 in Appendix A). illustrates the prompt example for ad generation tasks. To mitigate noise, the comparison is repeated $K$ times (typically an even number to evenly swap orders to remove positional bias (shi2024judging), and ties are broken by random. Winners advance round by round until only the final winner remains. Given the final winner’s index $i^{\star}$, we define $(\Pi^{j}_{t,g},c^{j}_{t,g}):=(\Pi^{j,(i^{\star})}_{t,g},c^{j,(i^{\star})}_{t,g})$ and carry it forward for the next gradient step $g+1$.

### v3

@@ 3 Problem Definition
    Evaluation: The system outcome $\Phi(\pi_{t})$ is deployed, and its performance is measured, yielding a score $s_{t}$. The new observation is added to the history: $H_{t}=H_{t-1}\cup\{(\pi_{t},\Phi\left(\pi_{t}\right),s_{t})\}$.
We will be more precise about how Steps 1 and 2 (analysis and generation) are implemented in our self-improving AI framework in $\S$5 and $\S$6.
Optimization objective:
In this paper, we formulate the goal of self-improving AI algorithms $\mathcal{A}$ as improving prompts $\left(\pi_{t}\right)_{t=1}^{T}$ over the evaluation budget $T$ to efficiently optimize the objective $J(\pi)$. Specifically, we discuss evaluation efficiency using two standard regret criteria. Let $\pi^{\star}\in\arg\max_{\pi\in\Pi}J(\pi)$. Then,
  • $\bullet$
    The (expected) cumulative regret is $\mathbb{E}\left[\sum_{t=1}^{T}\left(J\left(\pi^{\star}\right)-J\left(\pi_{t}\right)\right)\right]$.
  • $\bullet$
    The (expected) simple regret is $\mathbb{E}\left[J\left(\pi^{\star}\right)-\max_{1\leq t\leq T}J\left(\pi_{t}\right)\right]$.
In the next section, we show how a self-improving AI algorithm can be viewed as UCB-based Bayesian Optimization, which is optimal under both regret criteria (whitehouse2023sublinear; wang2023regret).
4 Our Approach: Recasting Iterative Prompt Optimization as Bayesian Optimization in Language Space
As discussed in $\S$1, our key idea is to model prompt-based self-improving AI as a Bayesian Optimization (BO) problem in language space. At a high level, we want a self-improving AI algorithm that repeatedly chooses a prompt, observes a noisy scalar performance signal (e.g., ad effectiveness), and decides which prompt to try next under a limited evaluation budget. We note that our optimization problem has no closed-form expression mapping the prompt to a performance score and a key constraint is that evaluating the function is costly. This is precisely the regime where BO has been shown to be effective, i.e., black-box objective functions where function evaluation is costly (frazier2018tutorial; srinivas2012information; papenmeier2025understanding).
Canonical UCB Bayesian Optimization.
To fix ideas, we start with a quick overview of BO. Classical BO considers a black-box function $f:\mathcal{Z}\to\mathbb{R}$ defined over a continuous input space $\mathcal{Z}$. The algorithm maintains a surrogate model—typically a Gaussian Process or related probabilistic regressor—that, given a history of evaluations $H_{t-1}=\{(z_{\tau},y_{\tau})\}_{\tau=1}^{t-1}$, produces for each $z\in\mathcal{Z}$ a posterior mean $\mu_{t-1}(z)$ and an uncertainty estimate $\sigma_{t-1}(z)$. Rather than maximizing $\mu_{t-1}(z)$ directly, BO selects the next query by maximizing an acquisition function that trades off exploitation (high predicted value) and exploration (high uncertainty). A widely used acquisition is the Upper Confidence Bound (UCB)
$A_{\beta_{t}}(z)\;=\;\mu_{t-1}(z)\;+\;\beta_{t}\,\sigma_{t-1}(z),$ (5)
where $\beta_{t}>0$ controls the strength of exploration (srinivas2012information). Figure 2 illustrates how $A_{\beta_{t}}$ combines the surrogate mean and uncertainty. Intuitively, $A_{\beta_{t}}(z)$ implements an “optimism in the face of uncertainty” rule: points with large uncertainty receive a positive bonus even if their current mean is modest.
Figure 2: Illustration of the upper confidence bound (UCB) acquisition function in Bayesian optimization. The unknown objective $f(z)$ (orange) is evaluated at a set of observed points ($\times$). A surrogate model yields the posterior mean $\mu(z)$ (green dashed) and $\sigma(z)$. The UCB acquisition $A_{\beta}(z)=\mu(z)+\beta\sigma(z)$ (blue) trades off exploitation and exploration by favoring locations with high predicted value and high uncertainty.
Input: Evaluation budget $T$, exploration parameter schedule $\beta$, initial history $H_{0}=\emptyset$
Output: Sequence of evaluation points $\{z_{t}\}_{t=1}^{T}$ and observations $\{y_{t}\}_{t=1}^{T}$
1 for $t=1,\dots,T$ do
2   Update surrogate model $p(f\mid H_{t-1})$ on past data
3   Construct the UCB acquisition
$A_{\beta}(z)\;=\;\mu_{t-1}(z)\;+\;\beta\,\sigma_{t-1}(z).$
Select the next evaluation point
$z_{t}\;\in\;\operatorname{argmax}_{z}A_{\beta}(z).$
Evaluate $f$ at $z_{t}$ to obtain a noisy observation
$y_{t}=f(z_{t})+\text{noise}.$
Augment the data: $H_{t}\leftarrow H_{t-1}\cup\{(z_{t},y_{t})\}$
4
Algorithm 1 Canonical UCB Bayesian Optimization
Algorithm 1 summarizes the resulting UCB-BO loop. At each iteration $t$, the surrogate is updated on $H_{t-1}$, the acquisition $A_{\beta_{t}}$ is constructed, the next point $z_{t}$ is chosen by maximizing $A_{\beta_{t}}$, the black-box function $f$ is evaluated at $z_{t}$ to obtain $y_{t}$, and the history is augmented. Under mild conditions, UCB-BO achieves sublinear regret and enjoys strong evaluation-efficiency guarantees: no algorithm can perform uniformly better in terms of regret scaling (srinivas2012information; whitehouse2023sublinear; wang2023regret).
In many practical settings, the global maximizer of $A_{\beta_{t}}$ is not available in closed form, and the acquisition is highly non-convex. A standard remedy is parallel gradient-based BO (wilson2018maximizing; papenmeier2025understanding): one starts from multiple initial points, performs gradient ascent on $A_{\beta_{t}}$ from each initialization to find local maxima, and then evaluates $f$ at these local optima in parallel. This procedure has shown surprising empirical effectiveness in high-dimensional numerical optimization (papenmeier2025understanding).
Mapping self-improving AI to UCB-BO.
We now conceptually align our iterative prompt-optimization problem with the BO template above. The correspondence is:
  • $\bullet$
    The BO input space $\mathcal{Z}$ corresponds to the prompt space $\Pi$.
  • $\bullet$
    The black-box function $f(z)$ corresponds to the system-level objective $f(\pi)\equiv J(\pi):=\mathbb{E}[r(y(\Phi(\pi),x))]$, where $x\sim\mathcal{D}_{\mathcal{X}}(\pi)$ and $r$ is the scoring rule defined in § 3.
  • $\bullet$
    The evaluation of $f(\pi)$ is the costly step: it requires deploying $\Phi(\pi)$, interacting with the environment (e.g., ad platform users), and computing $s_{t}$ as an empirical estimate of $J(\pi)$.
  • $\bullet$
    The cheap computation is everything the LLM-based system can do without new evaluations: analyzing past history, proposing textual edits, or internally comparing candidate ads.
If we had an explicit surrogate model over prompts, providing $\mu_{t}(\pi)$ and $\sigma_{t}(\pi)$ for each $\pi\in\Pi$, we could in principle construct a UCB acquisition
$A_{\beta_{t}}(\pi)\;=\;\mu_{t}(\pi)\;+\;\beta_{t}\,\sigma_{t}(\pi),$ (6)
and then run parallel gradient-based BO over $\Pi$: start from multiple initial prompts, ascend the gradient $\nabla_{\pi}A_{\beta_{t}}(\pi)$ to obtain local optima, and evaluate $J(\pi)$ only at those local optima. This would give us a self-improving AI algorithm that is, by construction, evaluation-efficient. However, directly instantiating this plan in language space runs into two central challenges (originally discussed in $\S$1).
Challenge 1: No native gradients in prompt space.
Prompts are discrete sequences of tokens, not elements of a Euclidean vector space. As such, prompt space lacks the inner-product structure of a Hilbert space that underlies standard notions of directions, norms, and gradients (young1988introduction). Expressions such as $\nabla_{\pi}A_{\beta_{t}}(\pi)$ are therefore not well-defined in the usual calculus sense.
A seemingly natural workaround is to embed prompts into a continuous space, take gradients there, and then “project back” to text:
$\pi\;\xrightarrow{\text{embed}}\;e=\mathsf{E}(\pi)\in\mathbb{R}^{d}\;\xrightarrow{\text{gradient step}}\;e^{\prime}\;\xrightarrow{\text{decode}}\;\pi^{\prime}=\mathsf{E}^{-1}(e^{\prime}).$ (7)
where $\mathsf{E}$ is the prompt embedding. Such a strategy is indeed theoretically plausible thanks to the invertibility of prompt embedding (nikolaou2025language). However, in practice, it encounters the well-known soft-to-discrete projection problem: learning the inverse prompt embedding ($\mathsf{E}^{-1}$), i.e., learning to decode an embedding into a prompt, is highly challenging (cui2025automatic). As a result, we cannot reliably implement gradient ascent in prompt space by differentiating through embeddings.
Challenge 2: No explicit UCB or uncertainty model.
Even if we had a notion of gradient in language space, we would still face a second obstacle: the UCB acquisition function itself is not explicitly available. In standard BO, the surrogate model exposes both a posterior mean $\mu_{t}(\cdot)$ and an uncertainty estimate $\sigma_{t}(\cdot)$, from which we build $A_{\beta_{t}}(\cdot)$ and compute its gradient.
In our setting, by contrast, there is no explicit Gaussian process or parametric surrogate over prompts. The relevant beliefs about prompt quality and uncertainty live implicitly inside the LLM and its critic: they are encoded in how the model generalizes from past evaluations, not in an externally accessible $\mu_{t}$ or $\sigma_{t}$. Consequently, we cannot simply write down $A_{\beta_{t}}(\pi)$, evaluate it at arbitrary prompts, or compute $\nabla_{\pi}A_{\beta_{t}}(\pi)$ in closed form. Naive proxies for uncertainty (e.g., self-consistency variance, embedding distances) tend to be poorly calibrated or brittle, and training a separate Gaussian Process over high-dimensional text embeddings is often expensive, data-hungry, and empirically unstable.
Textual gradients as a partial solution.
Recent work on textual gradients offers a way to tackle the first challenge. The TextGrad framework (textgrad) shows that one can approximate gradient ascent on an implicit objective by asking an auxiliary LLM critic to propose local natural-language edits that are predicted to improve a scalar score. Applying these edits to the current prompt yields a sequence of locally improving prompts, which empirically and theoretically behaves as gradient ascent on the posterior mean $\mu(\pi)$ in an implicit embedding space thanks to bi-Lipschitzness of LLM’s prompt embedding (nikolaou2025language; tang2025understanding).
However, textual gradients as in textgrad are inherently exploitative: they aim to climb the posterior mean $\mu(\pi)$, ignoring epistemic uncertainty $\sigma(\pi)$. They, therefore, correspond to gradient ascent on $\mu(\pi)$, not on the UCB acquisition $A_{\beta_{t}}(\pi)=\mu(\pi)+\beta_{t}\sigma(\pi)$. To recover the evaluation-efficiency guarantees of UCB-BO, we need an update rule whose effective search direction behaves like $\nabla_{\pi}A_{\beta_{t}}(\pi)$, even though we never observe $\sigma(\pi)$ or $\nabla_{\pi}\sigma(\pi)$ explicitly.
Our approach.
The central technical idea of this paper is to address Challenge 1 and Challenge 2 jointly by combining textual gradients with a Best-of-N selection mechanism. Rather than differentiating an explicit acquisition function, we (i) use a critic LLM to generate multiple stochastic textual edits around the current prompt, (ii) apply these edits to obtain nearby candidate prompts and system outcomes, and (iii) ask the critic, conditioned on past history, to pick the most promising candidate in a Best-of-N fashion.
In $\S$5, we formalize this procedure as a Best-of-N gradient and show that, under mild regularity conditions, the direction selected by this mechanism converges (in probability) to the ascent direction of a UCB-style acquisition function in the LLM’s implicit embedding space, with exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 1). This theoretical result gives us a principled way to emulate parallel gradient-based UCB-BO entirely in language space using only black-box text interactions. We then leverage this equivalence to design a concrete self-improving AI algorithm in $\S$6, TextBO, that is evaluation-efficient, simple to implement, and readily deployable in real-world applications such as digital advertising.
@@ 5 Best-of-N gradient and Theory
As discussed in $\S$4, our aim is to implement UCB-style Bayesian Optimization in prompt space using only black-box access to an LLM critic. The theoretical challenge is to show that the combination of (i) textual gradients and (ii) Best-of-N selection actually induces gradient ascent on a UCB acquisition function in the LLM’s implicit embedding space. This section formalizes that idea and proves our main result.
We proceed in three steps. In $\S$5.1, we formalize the role of the critic LLM, which provides the primitive operations—pairwise judgments, textual gradients, and meta-reflections—that enable search in language space. $\S$5.2 then introduces the Best-of-N gradient, our key mechanism for turning these critic capabilities into a stochastic, locally improving update rule. Finally, in $\S$5.3, we prove that Best-of-N gradients asymptotically emulate gradient ascent on a UCB acquisition function in the LLM’s implicit embedding space. This establishes the theoretical justification for the TextBO framework developed in the next section.
5.1 Critic LLM
The use of a separate LLM as a critic is now a standard pattern in the prompt-optimization and self-improving AI literature: critic models are routinely used to provide structured feedback, propose improvements, or act as reward models and judges that guide search over prompts and outputs (zhou2022large; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro; shi2024judging). We therefore treat the critic LLM as a natural and practically realistic component that supplies the basic operations needed to implement Bayesian-optimization-style updates in language space.
Recall that we denote a prompt as $\pi$ and the corresponding candidate solution (e.g., an ad creative) as $\Phi(\pi)$, the optimization history up to iteration $t$ as $H_{t}=\{(\pi_{\tau},\Phi(\pi_{\tau}),s_{\tau})\}_{\tau\leq t}$, where $s_{\tau}$ is the evaluated score of $\pi_{\tau}$. Then, let $M_{\text{critic}}$ denote the critic LLM, which can perform the following four operations:
@@ 5.1 Critic LLM
    Pairwise Judge Operator. Given two prompts $\pi^{(1)}$ and $\pi^{(2)}$, and their corresponding outcomes $\Phi(\pi^{(1)})$ and $\Phi(\pi^{(2)})$, predict which is better given reflection $R$:
    $\textsc{PairwiseJudge}(\Phi(\pi^{(1)}),\Phi(\pi^{(2)});R,M_{\text{critic}})\in\{\Phi(\pi^{(1)}),\Phi(\pi^{(2)})\}.$ (8)
@@
    Meta-reflection Operator. Compress the entire history into a small set of natural-language rules or guidelines:
    $R_{t}=\textsc{MetaReflect}(H_{t};M_{\text{critic}}).$ (9)
@@
    Textual Gradient Operator. Stochastically generate a local edit $\delta$ intended to improve $\pi$ under reflection $R$:
    $\delta\sim\nabla_{\text{text}}(\pi;R,M_{\text{critic}}).$ (10)
@@
    Apply Operator. Apply an edit $\delta$ to obtain a new prompt:
    $\pi^{\prime}=\operatorname{Apply}(\pi,\delta;M_{\text{critic}}).$ (11)
These four operators map directly onto well-documented LLM capabilities in the prompt-optimization and evaluative-feedback literature. The PairwiseJudge operator aligns with the growing use of LLMs as reliable comparison judges for creative and reasoning tasks (shi2024judging; calderon2025alternative). The MetaReflect operator captures reflection-based prompt-improvement methods that distill past outcomes into actionable guidelines (e.g., in the ad optimization case, it could suggest that “social settings outperform isolated product shots") (jiang2024many; ferber2024context; wang2025maestro). The TextualGradient operator corresponds to LLM-driven proposals for targeted, locally improving edits (e.g., “make the lighting warmer and emphasize people enjoying the product”), as formalized in textgrad and applied across domains such as scientific reasoning and code refinement (li2025test). Finally, the Apply operator reflects standard LLM-based prompt editing widely used in prompt engineering (zhou2022large; khattab2023dspy). Together, these capabilities make an LLM critic a natural, practically grounded component for enabling BO-style search over prompt space.
@@ 5.2 Best-of-N gradient
The notion of a Best-of-N gradient builds on the widely used Best-of-N strategy for test-time alignment, where an LLM trades additional inference-time compute for higher-quality outputs by sampling multiple candidates and then selecting the one that appears most aligned with a downstream objective (snell2024scaling; beirami2024theoretical). In standard Best-of-N prompting, an LLM generates $N$ candidate prompts, a generation model produces $N$ candidate outcomes, and a reward model selects the highest-scoring one. While effective for choosing a single output, our goal here is to improve the prompt itself. To accomplish this, we introduce the concept of Best-of-N gradient. Rather than sampling $N$ full prompts, we sample $N$ distinct textual gradients, apply them to the current prompt, and then select the candidate whose predicted outcome is judged most promising. This selection defines a localized “direction of improvement” in prompt space, and can be viewed as an acquisition function.
Formally, given a current prompt $\pi$ and reflection $R$, a single Best-of-N gradient step proceeds as follows:
@@
    Textual gradient generation. Use $M_{\text{critic}}$ to generate $N$ independent textual edits:
    $\delta^{(i)}\sim\nabla_{\text{text}}(\pi;R,M_{\text{critic}}),\qquad i=1,\dots,N.$ (12)
@@
    Candidate prompt generation. Apply each edit to produce $N$ candidate prompts:
    $\pi^{(i)}=\operatorname{Apply}(\pi,\delta^{(i)};M_{\text{critic}}).$ (13)
@@
    Candidate system outcome generation. Produce the associated system outcomes:
    $\Phi(\pi^{(i)}),\qquad i=1,\dots,N.$ (14)
@@
    Best-of-N selection. Identify the most promising candidate using the critic:
    $i^{\star}=\textsc{Best-of-N}\big(\{(\pi^{(i)},\Phi(\pi^{(i)})\}_{i=1}^{N};R,M_{\text{critic}}\big).$ (15)
    For the Best-of-N selection, we employ a pairwise tournament (liu2025pairjudgermperformbestofn). The $N$ candidates are first randomly permuted into tournament buckets. Each match compares two outcomes using the PairwiseJudge operator, with the comparison repeated $K$ times and the order of presentation swapped to mitigate positional bias (shi2024judging) (see the prompt template in Figure A3 in Web Appendix C.3 for the ad-optimization case). Ties are broken at random. Winners advance round by round until a single final winner remains; its associated prompt and outcome are then carried forward to the next gradient step.
The resulting prompt $\pi^{(i^{\star})}$ defines the update made by the Best-of-N gradient for a given step. Intuitively, each textual edit represents a noisy proposal for a local improvement direction; Best-of-N filters these proposals through an optimism-based rule, selecting the direction most likely to improve performance under the critic LLM’s current understanding.